AIDAN SHARPE

(609) 738-5237 ⋄ Hightstown, NJ amsharpe102@pm.me ⋄ Linked In ⋄ GitHub

GROWTH DRIVEN ENGINEER • 8 years of diverse programming & electronics experience

• Practical skills in leading myself and others to accomplish goals

• Active and highly motivated learner with a creative and ambitious mindset

EDUCATION Bachelor of Electrical and Computer Engineering, Rowan University Expected May, 2025 GPA: 3.74 Martinson Honor’s College Fall 2021 – Present Dean’s List: Fall 2021, Spring 2022 Relevant Coursework: Circuit Design & Analysis, Computer Architecture, Embedded Systems, FPGA Programming

SKILLS

(5)* Programming  C, Python, C++, Verilog HDL, C# for .NET Xamarin, Java
(4) Embedded Systems  Experience programming MSP430 and Arduino devices
(4) IoT Systems  Microsoft Azure IoTHub, Docker, Podman, Ubuntu Server, Fedora Server
(4) Electronics  Circuit & PCB Design, Through-hole & Surface Mount Soldering
(4) Hardware Testing  Examines circuits with Waveform Generators, Multimeters, & Oscilloscopes
(4) Robotics  Subsystem command programming, Autonomous behavior, Data & Power Cabling
(5) General Software  Fedora, NixOS, Debian-based Linux, Windows 10/11, Microsoft Office, Google Office
(5) Special Software  OnShape, Wireshark, SPICE, GitHub, VS Code, Vim, VMWare, VirtualBox
(5) Soft Skills  Professional Written & Verbal Communication, Leadership, Initiative

*Skill level rating; 5 highest

WORK EXPERIENCE Laser Fleet Sailing Coach Jun 2020 - August 2022 Toms River Yacht Club Toms River, NJ

• Greatly advanced the abilities and confidence of beginner Laser sailors

• Orchestrated and presented lessons to boost the skills of students Technology Development Engineering Intern May 2023 - August 2023 Inductotherm Corp. Rancocas, NJ

• Improved reliability and upgradability by automating remote configuration processes

• Enabled more efficient deployments of gateway computers by using NixOS

• Improved operation of test fixture using a custom designed PCBA

AWARDS & DISTINCTIONS • Eagle Scout with Bronze & Gold Palms 2020 • Smart City Innovation Track Winner at 24-hour rapid engineering competition 2022

– Innovated communication with an open-source digital assistant

• National Honor Society & German National Honor Society Inducted 2020 • Professional Engineering Society of Mercer County — Engineering Student Scholarship 2021 • FIRST Robotics Competition Team Mercury: 1089 Gracious Professionalism Scholarship 2021 EXTRA-CURRICULAR ACTIVITIES

• Actively document & pioneer proof-of-concept programming projects. 2017 – Present • Rowan University Intramural Flag Football to balance work and an active lifestyle Fall 2022 • Rowan IEEE Student Branch Engineering Workshop Leader 2021 – Present • Rowan Circle K International 2022 – Present • FIRST Robotics Competition 2017 – 2021